# Dirac Sources for Nonmetricity and Torsion in Metric-affine Gravity

### v1

@@ Abstract
Metric-affine gravity (GL(4) gauge theory) in 4-dimensions is coupled to a spacetime Dirac source field using the isomorphisms of the Lie algebra gl(4) to the Clifford algebras Cl(3,1) and Cl(2,2). A simple transformation relates the generators of Cl(3,1) to a real representation of Cl(2,2), while the real representation of Cl(2,2) serves directly as a basis for the Lie algebra gl(4). Therefore, although GL(4) does not contain a spinor representation of the Lorentz group, expanding its Lie algebra in the Cl(2,2) basis gives a Clifford valued connection with well-defined coupling to Dirac spinors. Variation of the expansion coefficients gives new Dirac sources for both torsion and nonmetricity, separated by identifying the so(3,1) basis within the gl(4) basis.
@@ 1 Introduction
so that $\mathfrak{T}_{\;\;\;\mu\nu}^{a}=\Sigma_{\;\;\;\nu\mu}^{a}-\Sigma_{\;\;\;\mu\nu}^{a}$. Here we use Latin indices for orthonormal frames and Greek indices for coordinate frames.
@@
The goal for our current work is to identify Dirac sources for both torsion and nonmetricity in $GL\left(4\right)$ gravity. It is important to observe that since the sources for torsion and nonmetricity both depend on the overall symmetry of the connection, the source for torsion differs between ECSK and $GL\left(4\right)$ (metric-affine) gravity. Concretely, Poincaré gauge theory couples only the totally antisymmetric part of the torsion to a single Dirac pseudo-current, but as we show here, all 16 Dirac currents play a role in metric-affine gravity. Since $GL\left(4\right)$ does not have a preferred metric, the sources for torsion differ strongly from the predictions of ECSK theory. The full range of Dirac sources for generic nonmetricity have not previously been identified.
@@
Since the Yang-Mills and scalar fields of the standard model do not provide sources for nonmetricity, excluding spinor sources ignores direct coupling of nonmetricity to the Standard Model. Nonetheless, coupling is possible because the connection depends only on the Lie algebra $\mathfrak{gl}\left(4\right)$, which does have spinor representations. An interesting recent study [25] includes Dirac sources, but unlike our presentation restricts the fermion coupling to the antisymmetric part of the connection. This restriction means that [25] only couples spinors to traces of the torsion and nonmetricity. Our current work couples the full connection to spinors.
@@
A second seeming obstacle to finding spinor sources in $GL\left(4\right)$ gravity is the definition of the spinors themselves, because spinors are defined as representations of a Clifford algebra which in turn depends on the Lorentz metric $\eta_{ab}$ through the definition of the gamma matrices, $\left\{\gamma^{a},\gamma^{b}\right\}=-2\eta^{ab}1$. The resolution lies in the independence of metric and connection in metric-affine gravity, which allows us to choose the Minkowski metric. Since any metric we choose will be not be compatible with the connection, we have this freedom.
@@
$e\Delta_{i}^{\;\;\;jk}\equiv-\frac{\delta\mathcal{L}_{source}}{\delta\Sigma_{\;\;\;jk}^{i}}$
where the metric and connection are independent. The source tensor $\Delta_{i}^{\;\;\;jk}$ is called the hypermomentum. By expanding the $GL\left(4\right)$ connection $\Sigma_{\;\;\;jk}^{i}$ in a $\mathfrak{cl}\left(3,1\right)$ basis we can describe the hypermomentum of the spinor field. The process is simplified by working with the Clifford algebra $\mathfrak{Cl}\left(2,2\right)$ built from the basis gamma matrices of $\mathfrak{spin}\left(2,2\right)$, which has real representations. Since $\mathfrak{Cl}\left(2,2\right)$ has real representations, a real basis for $\mathfrak{Cl}\left(2,2\right)$ may serve directly as a basis for $\mathfrak{gl}\left(4\right)$. At the same time there is a simple transformation from that real basis to a basis for $\mathfrak{Cl}\left(3,1\right)$.
@@
In the next Section we expand the metric-affine curvature, separating the torsion and nonmetric dependence from the usual Riemannian curvature, then vary to find the gravitational contribution to the torsion and nonmetric field equations. We show the well-known result that in vacuum the torsion vanishes; for the nonmetric variation we find that only the totally symmetric part of the nonmetricity vanishes. The remaining part of the nonmetricity consists of the Weyl vector and an antisymmetric piece which may be absorbed into the torsion [26]. In Section 3 we present details of the Clifford algebras $\mathfrak{Cl}\left(3,1\right)$ and $\mathfrak{Cl}\left(2,2\right)$, then write the coupled form of the Dirac action. In Section 4 we separate the generators of the $SO\left(3,1\right)$ subgroup of $GL\left(4\right)$, allowing us to identify the part of the connection variation relating to torsion. Our main results, the Dirac sources for torsion and nonmetricity in metric-affine gravity, are presented in Section 5, with some simple cases in the short final Section.
For the orthonormal vector basis we use lower case Latin letters, $\eta_{ab},\mathbf{e}^{a}$. For coordinates, if needed, we use lower case Greek, $g_{\alpha\beta},v^{\mu}$. For spinors we use upper case Latin indices, e.g., $\psi^{A},\left[\gamma^{a}\right]_{\;\;\;C}^{B}$, and for the Clifford basis we use upper case Greek, $\Gamma^{\Delta}\in\left\{1,\gamma^{a},\sigma^{ab},\gamma_{5}\gamma^{a},\gamma_{5}\right\}$. The Clifford basis may be divided into symmetric $\Gamma^{\Delta_{s}}$ and antisymmetric $\Gamma^{\Delta_{a}}$ parts. When needed we use the Dirac form of the gamma matrices, given in 3.
@@ 2.1 Decomposition of the connection and curvature
The spacetime arena we consider is the class of geometries modeled on the principle $GL\left(4\right)$ fiber bundle based on the quotient of the affine group $A\left(4\right)=R^{4}\rtimes GL\left(4\right)$ by the general linear group, $\mathcal{M}_{0}^{4}=A\left(4\right)/GL\left(4\right)$. Generalizing the homogeneous quotient manifold $\mathcal{M}_{0}^{4}\rightarrow\mathcal{M}^{4}$ and the flat Maurer-Cartan connection to $\boldsymbol{\Sigma}_{\;\;\;b}^{a}$, we have the Cartan structure equations:
@@
Choose the orthonormal basis, $g_{ab}=\eta_{ab}$, so that $\boldsymbol{\Sigma}_{ba}+\boldsymbol{\Sigma}_{ab}=-\mathbf{Q}_{ab}$ and let $\boldsymbol{\Omega}_{ab}=\boldsymbol{\Sigma}_{\left[ab\right]}$. Then the connection is
@@
where we define the nonmetric 2-form $\mathbf{Q}^{a}\equiv\frac{1}{2}\mathbf{e}^{b}\land\mathbf{Q}_{\;\;\;b}^{a}$. We may eliminate explicit dependence on the nonmetric 2-form $\mathbf{Q}^{a}$ by the field redefinition, $\mathbf{T}^{a}\equiv\boldsymbol{\mathfrak{T}}^{a}-\mathbf{Q}^{a}$, leaving the nonmetricity determined by its remaining, totally symmetric part $Q_{\left(abc\right)}$. The resulting structure equation determines a Lorentz connection with torsion
@@ 2.2 The action
In vacuum, $\mathscr{T}_{\;\;\;bc}^{a}=0$ and $\mathcal{\mathscr{Q}}_{cba}=0$. The torsion therefore vanishes, while contraction of $\mathcal{\mathscr{Q}}_{cba}=0$ shows that $Q_{\;\;\;ae}^{e}=\frac{1}{4}Q_{\;\;\;ea}^{e}=2W_{a}$ where $W_{b}$ is the Weyl vector. It follows that $Q_{c\left(ab\right)}=2\eta_{c(a}W_{b)}$ so that the full solution for the nonmetricity in vacuum is $Q_{abc}=2\eta_{c(a}W_{b)}+Q_{a\left[bc\right]}$ and we may write $\mathbf{Q}_{ab}$ in terms of $\boldsymbol{\omega}$ and $\mathbf{Q}^{a}$ only, both of which may be absorbed into a Weyl geometry with torsion [26].
@@ 3 Clifford basis and the metric-affine interaction
In nonflat spacetimes, the Dirac equation requires a covariant derivative. The connection becomes a Lie-algebra valued 1-form $\boldsymbol{\beta}_{A}G^{A}$ where generators $G^{A}$ form a basis for a spinor representation of the Lie algebra of the gravitational symmetry. In spaces with compatible $SO\left(p,q\right)$ invariant metric $\eta_{ab}$, this takes the form $\frac{1}{2}\boldsymbol{\omega}_{ab}\sigma^{ab}$ where the coefficients are characterized by antisymmetry $\boldsymbol{\omega}_{ab}=-\boldsymbol{\omega}_{ba}$. The $\left(p,q\right)$ signature then enters when an index is raised, $\boldsymbol{\omega}_{\;\;\;b}^{a}=\eta^{ac}\boldsymbol{\omega}_{cb}$. In particular this applies to a Lorentz connection, precluding any nonmetricity $\mathbf{Q}_{ab}$ since in an orthonormal basis nonmetricity depends entirely on the symmetric part $\mathbf{Q}_{ab}=-\boldsymbol{\omega}_{ab}-\boldsymbol{\omega}_{ba}$. Therefore, nonmetricity is exactly the Lorentz- or $SO\left(p,q\right)$-violating part of the spacetime connection.
@@ 3.1 Cl(3,1)
To make the couplings of Dirac fields to nonmetricity explicit, we use the Dirac representation, $\gamma^{a}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$, and $\gamma^{i}=\left(\begin{array}[]{cc}0&\sigma^{i}\\ -\sigma^{i}&0\end{array}\right)$. In this representation $\left[h\right]_{AB}=\left[\gamma^{0}\right]_{\;\;\;B}^{A}$, so the Hermitian inner product $\left\langle\chi,\psi\right\rangle=\chi^{\dagger}h\psi\equiv\bar{\chi}\psi$ we have $\bar{\chi}=\chi^{\dagger}h=\chi^{\dagger}\gamma^{0}$ in the usual way. In this representation, the remainder of the Clifford basis $\Gamma^{A}=\left\{1,\gamma^{a},\sigma^{ab},\gamma_{5}\gamma^{a},\gamma_{5}\right\}$ is
@@ 3.2 Cl(2,2)
Unlike $\mathfrak{spin}\left(3,1\right)$, the Lie algebra $\mathfrak{spin}\left(2,2\right)$ admits a real representation. This has the advantage of directly providing a basis for $\mathfrak{gl}\left(4,\mathbb{R}\right)$, while making only a slight change to the $\mathfrak{spin}\left(3,1\right)$ basis $\Gamma^{\Delta}$. As above, lower case Latin indices $a,b,\ldots=0,1,2,3,$ refer to either $SO\left(3,1\right)$ or $SO\left(2,2\right)$, while upper case Latin are spinor indices. It is convenient (but unusual!) to let $i,j,\ldots=0,1,3$.
@@ 3.3 The covariant Dirac equation
The spinor action must be made manifestly real, so we separate the real vacuum terms 14 from the interaction, $S_{D}=S_{D,V}-\alpha Re\int b_{aA}\psi^{\dagger}ih\gamma^{a}\hat{\Gamma}^{A}\psi$ to identify the interaction
@@
$\eta^{AC}\left[h\hat{\Gamma}\right]_{\left(CB\right)}$ $\in$ $\hat{\Gamma}^{\Delta_{s}}\left\{h1,h\gamma^{0},ih\sigma^{12},ih\sigma^{23},h\gamma_{5}\gamma^{1},h\gamma_{5}\gamma^{3},h\gamma^{1}h\gamma^{3},ih\sigma^{02},h\gamma_{5}\gamma^{0}\right\}$
are symmetric, giving generators
@@
Varying $\left(b_{cij},b_{c2},b_{c52},b_{c5}\right)_{i=0,1,3}$ will give the source for torsion while varying $\left(b_{c},b_{ci},b_{c2i},b_{c5i}\right)_{i=0,1,3}$ will give the source for nonmetricity.
@@ 5.2 Spinor interaction
For example, the third term in Eq.(34), $\alpha b_{31}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma^{1}\psi\right)$ becomes
@@ 5.3 Field equations
This is the $\mathfrak{spin}\left(2,2\right)$ decomposition of the torsion, with each independent projection sourced by a different current. The $\mathfrak{spin}\left(2,2\right)$ decomposition of the nonmetricity is comprised of the remaining 10 equations.
@@
$\mathcal{T}_{\;\;\;ab}^{0}$ $=$ $\frac{\alpha}{\kappa}\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&-I_{\bar{\nu}\rho}&-I_{\bar{\nu}\sigma}\\ 0&I_{\bar{\nu}\rho}&0&-I_{\bar{\rho}\sigma}\\ 0&I_{\bar{\nu}\sigma}&I_{\bar{\rho}\sigma}&0\end{array}\right)$
$\mathcal{T}_{\;\;\;ab}^{1}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&I_{\bar{\nu}\sigma}+I_{\bar{\mu}\rho}&I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}&0\\ -I_{\bar{\nu}\sigma}-I_{\bar{\mu}\rho}&0&0&I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}\\ -I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&0&0&I_{\bar{\mu}\rho}-I_{\bar{\nu}\sigma}\\ 0&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&I_{\bar{\nu}\sigma}-I_{\bar{\mu}\rho}&0\end{array}\right)$
$\mathcal{T}_{\;\;\;ab}^{2}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&R_{\bar{\mu}\rho}-R_{\bar{\nu}\sigma}&-R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}&\left(D_{\bar{\mu}\mu}-D_{\bar{\sigma}\sigma}\right)\\ R_{\bar{\nu}\sigma}-R_{\bar{\mu}\rho}&0&\left(D_{\bar{\nu}\nu}+D_{\bar{\rho}\rho}\right)&R_{\bar{\rho}\sigma}-R_{\bar{\mu}\nu}\\ R_{\bar{\mu}\nu}+R_{\bar{\rho}\sigma}&-\left(D_{\bar{\nu}\nu}+D_{\bar{\rho}\rho}\right)&0&-\left(R_{\bar{\mu}\rho}+R_{\bar{\nu}\sigma}\right)\\ -\left(D_{\bar{\mu}\mu}-D_{\bar{\sigma}\sigma}\right)&R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}&\left(R_{\bar{\mu}\rho}+R_{\bar{\nu}\sigma}\right)&0\end{array}\right)$
$\mathcal{T}_{\;\;\;ab}^{3}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&I_{\bar{\nu}\rho}-I_{\bar{\mu}\sigma}&0&-\left(I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}\right)\\ I_{\bar{\mu}\sigma}-I_{\bar{\nu}\rho}&0&I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}&0\\ 0&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&0&-\left(I_{\bar{\mu}\sigma}+I_{\bar{\nu}\rho}\right)\\ I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}&0&I_{\bar{\mu}\sigma}+I_{\bar{\nu}\rho}&0\end{array}\right)$

### v2

@@ Abstract
[…]
@@ 1 Introduction
so that $\mathfrak{T}_{\;\;\;\mu\nu}^{a}=\Sigma_{\;\;\;\nu\mu}^{a}-\Sigma_{\;\;\;\mu\nu}^{a}$. Here we use Latin indices $\left(a,b\ldots=0,1,2,3\right)$ for orthonormal frames and Greek indices $\left(\mu,\nu\ldots=0,1,2,3\right)$ for coordinate frames.
@@
The goal for our current work is to identify Dirac sources for both torsion and nonmetricity in $GL\left(4\right)$ gravity. It is important to observe that since the sources for torsion and nonmetricity both depend on the overall symmetry of the connection, the source for torsion differs between ECSK and $GL\left(4\right)$ (metric-affine) gravity. Concretely, Poincaré gauge theory couples only the totally antisymmetric part of the torsion to a single Dirac pseudo-current, but as we show here, all 16 Dirac currents play a role in metric-affine gravity. The $GL\left(4\right)$ sources for torsion differ strongly from the predictions of ECSK theory, while the full range of Dirac sources for generic nonmetricity have not previously been identified.
@@
Since the Yang-Mills and scalar fields of the standard model do not provide sources for nonmetricity, excluding spinor sources ignores direct coupling of nonmetricity to the Standard Model. Nonetheless, coupling is possible because the connection depends only on the Lie algebra $\mathfrak{gl}\left(4\right)$, which does have spinor representations. An interesting recent study [25] includes Dirac sources, but unlike our presentation [25] restricts the fermion coupling to the antisymmetric part of the connection. This restriction means that [25] only couples spinors to traces of the torsion and nonmetricity. Our current work couples the full connection to spinors.
@@
A second seeming obstacle to finding spinor sources in $GL\left(4\right)$ gravity is the definition of the spinors themselves, because spinors are defined as representations of a Clifford algebra. The Clifford algebra, in turn, depends on the Lorentz metric $\eta_{ab}$ through the definition of the gamma matrices, $\left\{\gamma^{a},\gamma^{b}\right\}=-2\eta^{ab}1$. The resolution lies in the independence of metric and connection in metric-affine gravity, which allows us to choose the Minkowski metric. Since any metric we choose will be not be compatible with the connection, we have this freedom.
@@
$e\Delta_{a}^{\;\;\;bc}\equiv-\frac{\delta\mathcal{L}_{source}}{\delta\Sigma_{\;\;\;bc}^{a}}$
where the metric and connection are independent. The source tensor $\Delta_{a}^{\;\;\;bc}$ is called the hypermomentum. By expanding the $GL\left(4\right)$ connection $\Sigma_{\;\;\;bc}^{a}$ in a $\mathfrak{cl}\left(3,1\right)$ basis we can describe the hypermomentum of the spinor field. The process is simplified by working with the Clifford algebra $\mathfrak{Cl}\left(2,2\right)$ built from a real basis of gamma matrices for $\mathfrak{spin}\left(2,2\right)$. The resulting real basis for $\mathfrak{Cl}\left(2,2\right)$ serves directly as a basis for $\mathfrak{gl}\left(4\right)$. At the same time there is a simple transformation from that real basis to a basis for $\mathfrak{Cl}\left(3,1\right)$.
@@
In Section 2 we expand the metric-affine curvature, separating the torsion and nonmetric dependence from the usual Riemannian curvature, then vary to find the gravitational contribution to the torsion and nonmetric field equations. We show the well-known result that in vacuum the torsion vanishes. The nonmetric variation shows that only the totally symmetric part of the nonmetricity vanishes. The remaining part of the nonmetricity consists of the Weyl vector and an antisymmetric piece which may be absorbed into the torsion [26]. In Section 3 we present details of the Clifford algebras $\mathfrak{Cl}\left(3,1\right)$ and $\mathfrak{Cl}\left(2,2\right)$, then write the coupled form of the Dirac action. In Section 4 we separate the generators of the $SO\left(3,1\right)$ subgroup of $GL\left(4\right)$, allowing us to identify the part of the connection variation relating to torsion. Our main results, the Dirac sources for torsion and nonmetricity in metric-affine gravity, are presented in Section 5, with a summary and some simple cases in the final Section.
For the orthonormal vector basis we use lower case Latin letters from the beginning of the alphabet, $\eta_{ab},\mathbf{e}^{a}\left(a,b,\ldots=0,1,2,3\right)$, with indices $i,j,\ldots=1,2,3$. For coordinates we use lower case Greek, $g_{\alpha\beta},v^{\mu}$. For spinors we use upper case Latin indices, e.g., $\psi^{A},\left[\gamma^{a}\right]_{\;\;\;C}^{B}$, and for the Clifford basis we use upper case Greek, $\Gamma^{\Delta}\in\left\{1,\gamma^{a},\sigma^{ab},\gamma_{5}\gamma^{a},\gamma_{5}\right\},\left(\Delta,\Xi,\ldots=1,\ldots,16\right)$. The Clifford basis may be divided into symmetric $\Gamma^{\Delta_{s}}$ and antisymmetric $\Gamma^{\Delta_{a}}$ parts. When needed we use the Dirac form of the gamma matrices, given in 3.
@@ 2.1 Decomposition of the connection and curvature
The spacetime arena we consider is the class of geometries modeled on the principle $GL\left(4\right)$ fiber bundle based on the quotient of the affine group $A\left(4\right)=R^{4}\rtimes GL\left(4\right)$ by the general linear group, $\mathcal{M}_{0}^{4}=A\left(4\right)/GL\left(4\right)$. Generalizing the homogeneous quotient manifold $\mathcal{M}_{0}^{4}\rightarrow\mathcal{M}^{4}$ and the flat Maurer-Cartan connection to the curved connection $\boldsymbol{\Sigma}_{\;\;\;b}^{a}$, we have the Cartan structure equations:
@@
We next separate the contributions of torsion and nonmetricity to the curvature. Choose the orthonormal basis, $g_{ab}=\eta_{ab}$, so that $\boldsymbol{\Sigma}_{ba}+\boldsymbol{\Sigma}_{ab}=-\mathbf{Q}_{ab}$ and let $\boldsymbol{\Omega}_{ab}=\boldsymbol{\Sigma}_{\left[ab\right]}$. Then the connection is
@@
where we define the nonmetric 2-form $\mathbf{Q}^{a}\equiv\frac{1}{2}\mathbf{e}^{b}\land\mathbf{Q}_{\;\;\;b}^{a}$. We may eliminate explicit dependence on the nonmetric 2-form $\mathbf{Q}^{a}$ by the field redefinition, ,
$\mathbf{T}^{a}$ $\equiv$ $\boldsymbol{\mathfrak{T}}^{a}-\mathbf{Q}^{a}$ (6)
leaving the nonmetricity determined by its remaining, totally symmetric part $Q_{\left(abc\right)}$. The resulting structure equation determines a Lorentz connection with torsion
@@
Equations (6) and (8) now express the original curvature and torsion in terms of $\mathbf{R}^{ab},\mathbf{Q}^{ab}$ and $\mathbf{T}^{a}$.
@@ 2.2 The action
In vacuum, $\mathscr{T}_{\;\;\;bc}^{a}=0$ and $\mathcal{\mathscr{Q}}_{cba}=0$. The torsion therefore vanishes, while contraction of $\mathcal{\mathscr{Q}}_{cba}=0$ shows that $Q_{\;\;\;ae}^{e}=\frac{1}{4}Q_{\;\;\;ea}^{e}=2W_{a}$ where $W_{b}$ is the Weyl vector. It follows that $Q_{c\left(ab\right)}=2\eta_{c(a}W_{b)}$ so that the full solution for the nonmetricity in vacuum is $Q_{abc}=2\eta_{c(a}W_{b)}+Q_{a\left[bc\right]}$. The full nonmetricity $\mathbf{Q}_{ab}$ is now given in terms of $\boldsymbol{\omega}$ and $\mathbf{Q}^{a}$ only, both of which may be absorbed into a Weyl geometry with torsion [26].
@@ 3 Clifford basis and the metric-affine interaction
In nonflat spacetimes, the Dirac equation requires a covariant derivative. The connection becomes a Lie-algebra valued 1-form $\boldsymbol{\beta}_{A}G^{A}$ where generators $G^{A}$ form a basis for a spinor representation of the Lie algebra of the gravitational symmetry. In spaces with compatible $SO\left(p,q\right)$ invariant metric $\eta_{ab}$, this takes the form $\frac{1}{2}\boldsymbol{\omega}_{ab}\sigma^{ab}$ where the coefficients are characterized by antisymmetry $\boldsymbol{\omega}_{ab}=-\boldsymbol{\omega}_{ba}$. The $\left(p,q\right)$ signature then enters when an index is raised, $\boldsymbol{\omega}_{\;\;\;b}^{a}=\eta^{ac}\boldsymbol{\omega}_{cb}$. In particular this applies to a Lorentz connection, precluding any nonmetricity $\mathbf{Q}_{ab}$ since in this basis nonmetricity depends entirely on the symmetric part $\mathbf{Q}_{ab}=-\boldsymbol{\omega}_{ab}-\boldsymbol{\omega}_{ba}$. Therefore, nonmetricity is exactly the Lorentz- or $SO\left(p,q\right)$-violating part of the spacetime connection.
@@ 3.1 Cl(3,1)
To make the couplings of Dirac fields to nonmetricity explicit, we use the Dirac representation, $\gamma^{a}=\left(\begin{array}[]{cc}1&0\\ 0&-1\end{array}\right)$, and $\gamma^{i}=\left(\begin{array}[]{cc}0&\sigma^{i}\\ -\sigma^{i}&0\end{array}\right)$ $\left(i=1,2,3\right)$. In this representation $\left[h\right]_{AB}=\left[\gamma^{0}\right]_{\;\;\;B}^{A}$, so the Hermitian inner product $\left\langle\chi,\psi\right\rangle=\chi^{\dagger}h\psi\equiv\bar{\chi}\psi$ we have $\bar{\chi}=\chi^{\dagger}h=\chi^{\dagger}\gamma^{0}$ in the usual way. The Dirac representation puts the remainder of the Clifford basis $\Gamma^{A}=\left\{1,\gamma^{a},\sigma^{ab},\gamma_{5}\gamma^{a},\gamma_{5}\right\}$ in the form
@@ 3.2 Cl(2,2)
Unlike $\mathfrak{spin}\left(3,1\right)$, the Lie algebra $\mathfrak{spin}\left(2,2\right)$ admits a real representation. This has the advantage of directly providing a basis for $\mathfrak{gl}\left(4,\mathbb{R}\right)$, while making only a slight change to the $\mathfrak{spin}\left(3,1\right)$ basis $\Gamma^{\Delta}$. As above, lower case Latin indices from the beginning of the alphabet $a,b,\ldots=0,1,2,3,$ refer to either $SO\left(3,1\right)$ or $SO\left(2,2\right)$, while upper case Latin are spinor indices.
@@ 3.3 The covariant Dirac equation
The spinor action must be made manifestly real, so we separate the real vacuum terms 15 from the interaction, $S_{D}=S_{D,V}-\alpha Re\int b_{aA}\psi^{\dagger}ih\gamma^{a}\hat{\Gamma}^{A}\psi$ to identify the interaction
@@
$\eta^{AC}\left[h\hat{\Gamma}\right]_{\left(CB\right)}$ $\in$ $\hat{\Gamma}^{\Delta_{s}}=\left\{h1,h\gamma^{0},ih\sigma^{12},ih\sigma^{23},h\gamma_{5}\gamma^{1},h\gamma_{5}\gamma^{3},h\gamma^{1}h\gamma^{3},ih\sigma^{02},h\gamma_{5}\gamma^{0}\right\}$
with $h\hat{\Gamma}$ symmetric, give generators
@@
Varying $\left(b_{c\mathfrak{ij}},b_{c2},b_{c52},b_{c5}\right)_{\mathfrak{i},\mathfrak{j}=0,1,3}$ will give the source for torsion while varying $\left(b_{c},b_{c\mathfrak{i}},b_{c2\mathfrak{i}},b_{c5\mathfrak{i}}\right)_{\mathfrak{i}=0,1,3}$ will give the source for nonmetricity.
@@ 5.2 Spinor interaction
where the over bar denotes complex conjugation. For example, the third term in Eq.(35), $\alpha b_{31}Re\left(\psi^{\dagger}ih\gamma^{3}\eta h\gamma^{1}\psi\right)$ becomes
@@ 5.3 Field equations
This is the $\mathfrak{spin}\left(2,2\right)$ decomposition of the torsion, with each independent projection sourced by a different current. The $\mathfrak{spin}\left(2,2\right)$ decomposition of the nonmetricity is comprised of the remaining 40 equations.
@@
$\mathcal{T}_{\;\;\;ab}^{0}$ $=$ $\frac{\alpha}{\kappa}\left(\begin{array}[]{cccc}0&0&0&0\\ 0&0&-I_{\bar{\nu}\rho}&-I_{\bar{\nu}\sigma}\\ 0&I_{\bar{\nu}\rho}&0&-I_{\bar{\rho}\sigma}\\ 0&I_{\bar{\nu}\sigma}&I_{\bar{\rho}\sigma}&0\end{array}\right)$ (42)
$\mathcal{T}_{\;\;\;ab}^{1}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&I_{\bar{\nu}\sigma}+I_{\bar{\mu}\rho}&I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}&0\\ -I_{\bar{\nu}\sigma}-I_{\bar{\mu}\rho}&0&0&I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}\\ -I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&0&0&I_{\bar{\mu}\rho}-I_{\bar{\nu}\sigma}\\ 0&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&I_{\bar{\nu}\sigma}-I_{\bar{\mu}\rho}&0\end{array}\right)$ (47)
$\mathcal{T}_{\;\;\;ab}^{2}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&R_{\bar{\mu}\rho}-R_{\bar{\nu}\sigma}&-R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}&\left(D_{\bar{\mu}\mu}-D_{\bar{\sigma}\sigma}\right)\\ R_{\bar{\nu}\sigma}-R_{\bar{\mu}\rho}&0&\left(D_{\bar{\nu}\nu}+D_{\bar{\rho}\rho}\right)&R_{\bar{\rho}\sigma}-R_{\bar{\mu}\nu}\\ R_{\bar{\mu}\nu}+R_{\bar{\rho}\sigma}&-\left(D_{\bar{\nu}\nu}+D_{\bar{\rho}\rho}\right)&0&-\left(R_{\bar{\mu}\rho}+R_{\bar{\nu}\sigma}\right)\\ -\left(D_{\bar{\mu}\mu}-D_{\bar{\sigma}\sigma}\right)&R_{\bar{\mu}\nu}-R_{\bar{\rho}\sigma}&\left(R_{\bar{\mu}\rho}+R_{\bar{\nu}\sigma}\right)&0\end{array}\right)$ (52)
$\mathcal{T}_{\;\;\;ab}^{3}$ $=$ $\frac{\alpha}{2\kappa}\left(\begin{array}[]{cccc}0&I_{\bar{\nu}\rho}-I_{\bar{\mu}\sigma}&0&-\left(I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}\right)\\ I_{\bar{\mu}\sigma}-I_{\bar{\nu}\rho}&0&I_{\bar{\mu}\nu}-I_{\bar{\rho}\sigma}&0\\ 0&I_{\bar{\rho}\sigma}-I_{\bar{\mu}\nu}&0&-\left(I_{\bar{\mu}\sigma}+I_{\bar{\nu}\rho}\right)\\ I_{\bar{\rho}\sigma}+I_{\bar{\mu}\nu}&0&I_{\bar{\mu}\sigma}+I_{\bar{\nu}\rho}&0\end{array}\right)$ (57)